OpenStudy (anonymous):

Use the elimination method to solve the given system of equations. Write your answer as an ordered pair (x,y). 5x -3y=8 2x +2y=16

OpenStudy (anonymous):

x=2,y=2/3

OpenStudy (anonymous):

Thanks Friend;)

myininaya (myininaya):

@scsam please provide the steps to getting the solution. It is against the Code of Conduct to just give the answer. Also @brandoduh you should be looking for more than just the answer. You should be looking for an explanation so you to can learn how to do these. Both parties need to engage in the learning process. The OpenStudy is not just an answer site. We want people to learn. Thanks and please. So if you wanted to do this using elimination We need to get the coefficients of either x to be opposite values or the coefficients of y to be opposite values So I think I will play with making the coefficients of y opposite values 5x -3y=8 2x +2y=16 So I know the lcm of 3 and 2 is 6 So I can multiply the first equation by 2 and the equation by 3 and I will get 10x-6y=16 6x+6y=48 ------------ Now adding we get 16x=64 solve this for x And then when you get this x value Plug into one of the equations given to find y

myininaya (myininaya):

And actually you won't get what @scsam got above

OpenStudy (anonymous):

@myininaya , @brandoduh I am so sorry about the answer. Here it is 5x-8=3y and 2x+2y=16 (5x-8)/3=y Substitute this in the other equation 2x+2*(5x-8)/3=16 6x+10x-16=48 x=4 Substitute the value of x in any of the above two equations, you get y=4 so x=4,y=4